# e-GPU: An Open-Source and Configurable RISC-V Graphic Processing Unit for TinyAI Applications

### v2

@@ Abstract
Graphics processing units (GPUs) excel at parallel processing, but remain largely unexplored in ultra-low-power edge devices (TinyAI) due to their power and area limitations, as well as the lack of suitable programming frameworks. To address these challenges, this work introduces embedded GPU (e-GPU), an open-source and configurable RISC-V GPU platform designed for TinyAI devices. Its extensive configurability enables area and power optimization, while a dedicated Tiny-OpenCL implementation provides a lightweight programming framework tailored to resource-constrained environments. To demonstrate its adaptability in real-world scenarios, we integrate the e-GPU with the eXtendible Heterogeneous Energy-Efficient Platform (X-HEEP) to realize an accelerated processing unit (APU) for TinyAI applications. Multiple instances of the proposed system, featuring varying e-GPU configurations, are implemented in TSMC’s SVT CMOS technology and are operated at and . Their area and leakage characteristics are analyzed to ensure alignment with TinyAI constraints. To assess both runtime overheads and application-level efficiency, we employ two benchmarks: General Matrix Multiply (GeMM) and bio-signal processing (TinyBio) workloads. The GeMM benchmark is used to quantify the scheduling overhead introduced by the Tiny-OpenCL framework. The results show that the delay becomes negligible for matrix sizes larger than (or equivalent problem sizes). The TinyBio benchmark is then used to evaluate performance and energy improvements in the baseline host. The results demonstrate that the high-range e-GPU configuration with 16 threads achieves up to a speed-up and reduces energy consumption by up to , while incurring only a area overhead and operating within a power budget.
Index Terms:
Ultra-Low-Power, Open-Source, Graphic Processing Unit, Microcontroller, Artificial Intelligence.I INTRODUCTION
@@
We release an open-source repository, including the complete e-GPU111The complete code and documentation of the e-GPU platform are freely available at https://github.com/esl-epfl/e-gpu. to allow researchers to tailor the platform to their TinyAI domains.
@@ VIII-C TinyBio Benchmark Characterization
Figure 4 presents the speed up and energy reduction achieved by each kernel of the TinyBio benchmark, described in Section VII-B, when executed on the analyzed e-GPU systems, relative to the baseline host system.
@@
Overall, when considering the complete application, the e-GPU systems achieve an overall performance improvement of to during the processing phases, along with an energy consumption reduction of to .
@@ IX CONCLUSION
The e-GPU is integrated with the X-HEEP host [10] to realize an APU optimized for TinyAI applications. Multiple instances of the proposed system, featuring varying e-GPU configurations, are implemented in TSMC’s SVT CMOS technology and are operated at and . Their area and leakage characteristics are analyzed to ensure alignment with TinyAI constraints. To assess both runtime overheads and application-level efficiency, we employed two benchmarks: GeMM and TinyBio workloads. The GeMM benchmark is used to quantify the scheduling overhead introduced by the Tiny-OpenCL framework. The results showed that the delay becomes negligible for matrix sizes larger than (or equivalent problem sizes). The TinyBio benchmark is then used to evaluate performance and energy improvements over the baseline host. The results demonstrated that the high-range e-GPU configuration achieves up to a speed-up and reduces energy consumption by up to , while incurring only a area overhead and operating within a power budget.

### v3

@@ Abstract
Graphics processing units (GPUs) excel at parallel processing, but remain largely unexplored in ultra-low-power edge devices (TinyAI) due to their power and area limitations, as well as the lack of suitable programming frameworks. To address these challenges, this work introduces embedded GPU (e-GPU), an open-source and configurable RISC-V GPU platform designed for TinyAI devices. Its extensive configurability enables area and power optimization, while a dedicated Tiny-OpenCL implementation provides a lightweight programming framework tailored to resource-constrained environments. To demonstrate its adaptability in real-world scenarios, we integrate the e-GPU with the eXtendible Heterogeneous Energy-Efficient Platform (X-HEEP) to realize an accelerated processing unit (APU) for TinyAI applications. Multiple instances of the proposed system, featuring varying e-GPU configurations, are implemented in TSMC’s SVT CMOS technology and are operated at and . Their area and leakage characteristics are analyzed to ensure alignment with TinyAI constraints. To assess both runtime overheads and computational efficiency, we employ two benchmarks: General Matrix Multiply (GeMM) and bio-signal processing (TinyBio) workloads. The GeMM benchmark is used to quantify the scheduling overhead introduced by the Tiny-OpenCL framework. The results show that the delay becomes negligible for matrix sizes larger than (or equivalent problem sizes). The TinyBio benchmark is then used to evaluate performance and energy improvements over the baseline host under pure processing conditions. The results indicate that the high-range e-GPU configuration with 16 threads achieves up to a speed-up and reduces energy consumption by up to , while incurring only a area overhead and operating within a power budget.
I INTRODUCTION
@@
We release an open-source repository, including the complete e-GPU111The complete code and documentation of the e-GPU platform is open-source on GitHub. to allow researchers to tailor the platform to their TinyAI domains.
@@ VIII-C TinyBio Benchmark Characterization
Figure 4 presents the speed up and energy reduction achieved by each kernel of the TinyBio benchmark, described in Section VII-B, when executed on the analyzed e-GPU systems, relative to the baseline host system. Since the impact of each overhead has been analyzed previously, transfer and scheduling overheads are excluded from these experiments, and only the pure computation phase is considered. Consequently, the data cache is assumed to be pre-populated with the input data required by each kernel, and a static kernel scheduling policy is employed.
@@
The SVM is executed after the feature extraction stage to perform final classification. Given its negligible execution time, it is included within the feature extraction phase.
Overall, the e-GPU systems achieve an average performance improvement ranging from to during the processing phases, along with an energy consumption reduction ranging from to .
@@ IX CONCLUSION
The e-GPU is integrated with the X-HEEP host [18] to realize an APU optimized for TinyAI applications. Multiple instances of the proposed system, featuring varying e-GPU configurations, are implemented in TSMC’s SVT CMOS technology and are operated at and . Their area and leakage characteristics are analyzed to ensure alignment with TinyAI constraints. To assess both runtime overheads and computational efficiency, we employ two benchmarks: GeMM and TinyBio workloads. The GeMM benchmark is used to quantify the scheduling overhead introduced by the Tiny-OpenCL framework. The results showed that the delay becomes negligible for matrix sizes larger than (or equivalent problem sizes). The TinyBio benchmark is then used to evaluate performance and energy improvements over the baseline host under pure processing conditions. The results indicate that the high-range e-GPU configuration with 16 threads achieves up to a speed-up and reduces energy consumption by up to , while incurring only a area overhead and operating within a power budget.